### v1

# Non-Gaussian state preparation and enhancement using weak-value amplification

Abstract
We introduce a protocol for generating a broad class of non-Gaussian (nG) quantum states via postselected weak measurement techniques. The scheme involves injecting an arbitrary quantum state and a single photon into the signal and idler ports, respectively, of an interference setup that incorporates a third-order nonlinear medium. A nG state is conditionally produced at the signal output, heralded by the detection of a single photon in one of the idler output channels. The protocol exploits a weak cross-Kerr interaction and effective single-photon nonlinearity enhanced by the weak-value amplification. We show that by tuning the weak value of the photon number operator in the idler mode within experimentally feasible parameters, a wide variety of nG states can be generated with high fidelity. As specific examples, we demonstrate the generation of photon-added states, displaced and squeezed number states, and a continuum of intermediate nG states using coherent and squeezed vacuum inputs, respectively. Furthermore, we show that the protocol enables the enhancement of non-Gaussianity and the enlargement of Schrödinger cat (SC) states when ideal SC states are used as the input. Our results provide an alternative route for the conditional generation of tunable nG states, with potential applications in quantum information processing. This approach may also open new avenues for quantum state engineering using postselected weak measurements.
I Introduction
Non-Gaussian (nG) states (Walschaers, 2021; Walschaers et al., 2020), defined as quantum states in the phase space that deviate from a Gaussian distribution, are often characterized by Wigner function negativity and have garnered significant interest owing to their enhanced utility across various domains of quantum physics. In particular, they play a central role in continuous-variable (CV) quantum information processing (Braunstein and van Loock, 2005), including quantum computation (Lloyd and Braunstein, 1999; Gottesman et al., 2001; Menicucci et al., 2006; Miyata et al., 2016; Mari and Eisert, 2012), quantum metrology (Anisimov et al., 2010; Huver et al., 2008), and foundational studies in quantum theory (Jeong et al., 2003; Stobińska et al., 2007). The generation of nG states requires either nG initial states or nG measurements; the latter alone is sufficient to produce non-Gaussianity. Among the existing techniques, nG unitary transformations and conditional measurements are the two most common approaches (Walschaers, 2021). Over the past decades, considerable efforts have focused on preparing nG states using conditional measurements involving nG quantum operations, such as photon addition and/or subtraction applied to Gaussian states in optical systems (Barbieri et al., 2010; Zavatta et al., 2004; Wakui et al., 2007; Neergaard-Nielsen et al., 2006; Wenger et al., 2004; Chen et al., 2024).
Squeezed single-photon and single-photon-added coherent (SPAC) states are among the most representative nG states and play irreplaceable roles in CV quantum information processing. These states can be prepared via nG operations such as single-photon addition and subtraction, applied to coherent and squeezed vacuum (SV) states, respectively. Another paradigmatic example is the Schrödinger cat (SC) state (Ourjoumtsev et al., 2006) inspired by Schrödinger’s famous thought experiment (Schrödinger, 1935). Current methods for generating SC states include photon-number-resolving detection (Dakna et al., 1997, 1999) and weak and strong nonlinear optical interactions (Yurke and Stoler, 1986; Gerry, 1999). Large-amplitude SC states characterized by coherent amplitudes $|\alpha|\geq 2$, where $\alpha$ is the complex amplitude, are particularly valuable for probing quantum foundations and enabling quantum information protocols (van Enk and Hirota, 2001; Ba An, 2003; An, 2004). However, SC states generated via photon subtraction or third-order optical nonlinearities typically have small amplitudes and fall short of the requirements imposed by advanced quantum tasks.
From a practical perspective, the purity of a state is critical for effective implementation of many quantum information protocols. However, due to limitations in photon detection efficiency and photon-number-resolving capabilities, generating high-fidelity nG states remains significantly more challenging than producing their Gaussian counterparts. This difficulty underscores the limitations of existing methods for preparing high-quality nG states. To address this challenge, one can either enhance the nGity of a given state through optimization techniques or develop alternative approaches capable of generating nG states with high purity and fidelity tailored for specific applications.
In this study, we propose an alternative protocol for generating a wide range of nG states based on the technique of weak-value amplification (WVA) (Aharonov et al., 1988; Dressel et al., 2014), without the need for explicit photon addition or subtraction operations. Our scheme relies on a weak cross-Kerr interaction between two optical modes, the signal and idler, mediated by third-order nonlinear susceptibility $\chi^{(3)}$. A nG state is conditionally generated in the signal output port from a given input signal state, heralded by single-photon detection in the idler output mode. The key element of the protocol is a postselected weak measurement of the photon number of the idler beam, which effectively induces single-photon nonlinearity via the WVA.
Although the protocol features a low probability of success, it enables the high-fidelity generation of various nG states. By appropriately tuning the weak value of the photon number operator in the idler mode within experimentally feasible parameters, our protocol can generate photon-added states and displaced number states from coherent input states. In the case of a squeezed vacuum (SV) input, the scheme yields high-fidelity two-photon-added SV states, squeezed number states, and equivalent squeezed even Schrödinger cat (SC) states. When applied to SC input states, the protocol serves two distinct purposes: (i) it significantly enhances the non-Gaussianity of the initial state, especially for small-amplitude even SC states, and (ii) it enables the expansion of small SC states into large-amplitude ($|\alpha|\geq 2$) SC states while preserving high fidelity ($F>0.99$). Additionally, our approach supports the generation of a continuous family of nonclassical states with intermediate properties corresponding to different classes of input states.
The remainder of this paper is organized as follows. In Sec. II, we present our nG state generation protocol and briefly discuss its feasibility for implementation in an optical laboratory. In Secs. III and IV, we provide a detailed theoretical analysis for generating the representative nG states. We demonstrate that for coherent and SV input states, our protocol can produce high-fidelity nonclassical states, including the SPAC state, displaced single-photon state, two-photon-added squeezed vacuum (TPASV) state, squeezed photon state, and squeezed even SC state. These states are of particular interest because of their potential applications in quantum information processing. For SC input states, we show that the protocol not only enhances the non-Gaussianity of the initial state, but also enables the preparation of large-amplitude SC states with $\alpha\geq 2$ and fidelity exceeding 0.99. Finally, we conclude the paper in Sec. V. Throughout this study, we adopted natural units with $\hbar=1$. All the numerical simulations were performed using the Python package QuTiP (Johansson et al., 2012, 2013).
II Model setup
A previous study (Hu et al., 2022) introduced a quantum state engineering scheme based on postselected weak measurements implemented in a medium characterized by the second-order nonlinear susceptibility of a BBO crystal. The associated interaction Hamiltonian is given by $H_{I}=\xi(a^{\dagger}b^{\dagger}-ab)$, where $a$ ($a^{\dagger}$) and $b$ ($b^{\dagger}$) are the annihilation (creation) operators of the signal and idler modes, respectively. In that proposal, the interaction strength $\chi^{(2)}$ is proportional to the second-order nonlinear susceptibility of the medium, that is $\xi\propto\chi^{(2)}$, and the signal and idle modes play the roles of the pointer and the measured system, respectively. In contrast, in many quantum optical protocols involving light-matter interactions, a different von Neumann-type interaction Hamiltonian is often employed, namely, the cross-Kerr interaction, described by
$H_{int}=\chi^{(3)}n_{a}n_{b},$ (1)
where $n_{a}=a^{\dagger}a$ and $n_{b}=b^{\dagger}b$ are the number operators for the signal and idler modes, respectively, and $\chi^{(3)}$ denotes the third-order nonlinear susceptibility of the medium. As demonstrated in previous work (Feizpour et al., 2011; Hallaji et al., 2017; Matsuoka et al., 2017), if the signal and idler beams are identified as the measured system and the pointer, respectively, the interaction Hamiltonian $H_{int}$ can be used within a von Neumann-type measurement framework. Analogously to the standard weak measurement scenario, in the Hamiltonian $H_{int}$, the photon number operator $n_{a}$ serves as the observable of the measured system, while $n_{b}$ acts as the pointer variable, canonically conjugate to the phase operator $\varphi$ of the radiation field. In this section, we demonstrate the utility of the interaction Hamiltonian $H_{int}$ for preparation of quantum states via a postselected weak measurement technique.
Figure 1: Schematic of the nG state generation protocol. In this scheme, a given state is prepared in the input signal mode, and a single-photon state is prepared in the idler mode, both of which pass through a Mach-Zehnder interferometer (MZI). Preselection is performed by sending the single photon through an unbalanced beam splitter (BS) with a small deviation $\epsilon$ from the ideal angle $\frac{\pi}{4}$. The signal and idler modes are then coupled via a Kerr medium. A single-photon detection event at one of the idler output ports (i.e., a click at detector $D_{1}$) heralds the desired output state in the signal mode.
The schematic setup of our state generation model, based on a postselected weak measurement characterized by the Hamiltonian interaction $H_{int}$, is shown in Fig. 1. As illustrated, the proposed setup consists of a Mach-Zehnder interferometer (MZI), Kerr medium, and several optical elements. In our model, the beam splitter (BS) plays a crucial role in the execution of the protocol. The action of the BS is described by the scattering matrix (Campos et al., 1989; Agarwal, 2013).
$U_{BS}=\left(\begin{array}[]{cc}\cos\theta e^{i\varphi_{\tau}}&\sin\theta e^{i\phi_{\rho}}\\ -\sin\theta e^{-i\phi_{\rho}}&\cos\theta e^{-i\varphi_{\tau}}\end{array}\right),$ (2)
where the reflectance and transmittance of the BS are given by $\sqrt{R}=\sin\theta e^{i\phi_{\rho}}$and $\sqrt{T}=\cos\theta e^{i\varphi_{\tau}}$, respectively, satisfying $R+T=1$. We assume that the signal mode is initially prepared in an arbitrary quantum state $|\phi\rangle$ and passes through the Kerr medium, while the idle mode is initialized in a single-photon state. As shown in Fig. 1, when the input to the first BS is $|1\rangle|0\rangle$ (single photon and vacuum), the resulting state after the BS is
$|\psi_{i}\rangle$ $=\cos\theta|0\rangle_{r}|1\rangle_{t}+i\sin\theta|1\rangle_{r}|0\rangle_{t}.$ (3)
assuming $\varphi_{\tau}=0$, and $\phi_{\rho}=\frac{\pi}{2}$. The interaction between the measured system and pointer occurs via the Kerr medium, described by the time-evolution operator:
[…]
where $g=\chi^{(3)}t$ is the interaction strength, with $t=L/\upsilon$ being the interaction time, $L$ is the length of the Kerr medium, and $\upsilon$ is the speed of light in the medium. Since $g$ is typically small, it suffices to expand the evolution operator to the first order as follows:
$U=\exp\left(-ig\hat{n}_{a}\otimes\hat{n}_{b}\right)\approx 1-ig\hat{n}_{a}\otimes\hat{n}_{b}$ (5)
Under this approximation, the initial state of the composite system evolves as
$|\Psi^{\prime}\rangle$ $=U|\psi_{i}\rangle\otimes|\phi\rangle=\exp\left(-ig\hat{n}_{a}\otimes\hat{n}_{b}\right)|\psi_{i}\rangle\otimes|\phi\rangle$
$\approx\left[1-ig\hat{n}_{a}\otimes\hat{n}_{b}\right]|\psi_{i}\rangle\otimes|\phi\rangle$ (6)
We now proceed to the postselection stage. After passing through a balanced ($50:50$) BS, we require that detector $D1$ registers a single photon while detector $D2$ remains inactive. This detection event corresponds to projecting onto the state:
[…]
Postselecting $|\psi_{f}\rangle$ from the state $|\Psi^{\prime}\rangle$, yields the (unnormalized) final pointer state:
$|\Phi\rangle$ $=\langle\psi_{f}|\psi_{i}\rangle\left[1-ig\langle n_{a}\rangle_{w}n_{b}\right]|\phi\rangle$
$=\langle\psi_{f}|\psi_{i}\rangle\left[|\phi\rangle-i\frac{g}{2}\left(1+\frac{1}{\epsilon}\right)b^{\dagger}b|\phi\rangle\right],$ (8)
where
$\langle n_{a}\rangle_{w}$ $=\frac{\langle\psi_{f}|a^{\dagger}a|\psi_{i}\rangle}{\langle\psi_{f}|\psi_{i}\rangle}=\frac{1}{2}+\frac{1}{2\epsilon}$ (9)
is the weak value of the photon number operator $n_{a}=a^{\dagger}a$ for the idler beam. This expression assumes a slight deviation from a $50:50$ BS, with $\theta=\frac{\pi}{4}-\epsilon$, where $\epsilon\ll 1$ is real. For a sufficiently small $\epsilon$, the weak value can exceed unity, as observed in previous studies (Feizpour et al., 2011; Hallaji et al., 2017). The success probability of the postselection (i.e., a click at $D1$ and no click at $D2$ ) is given by $|\langle\psi_{f}|\psi_{i}\rangle|^{2}=|\epsilon|^{2}$. The normalized output state of the signal mode is
$|\Phi_{f}\rangle=\frac{|\Phi\rangle}{\sqrt{p_{f}}},$ (10)
where $p_{f}=|\langle\Phi|\Phi\rangle|^{2}$ is the probability of success of the post-selection. From the form of $|\Phi_{f}\rangle$, it is evident that the output state is a superposition of the initial state $|\phi\rangle$ and the state $b^{\dagger}b|\phi\rangle$. The interference between these components enhances the quality of the generated state and can lead to non-trivial quantum effects. In particular, the term $b^{\dagger}b|\phi\rangle$ corresponds to a photon subtraction followed by photon addition, a process previously identified as a powerful tool for quantum state engineering (Parigi et al., 2007). Therefore, when the coefficient $\frac{g}{2}\left(1+\frac{1}{\epsilon}\right)\gg 1$, the $b^{\dagger}b|\phi\rangle$ component dominates the output, effectively transforming the initial state $|\phi\rangle$ into a distinct quantum state with nontrivial properties and nG characteristics.
As investigated in a recent WVA experiment (Li et al., 2018), the minimal rotation angle $\epsilon$ of the beam splitter (BS) can be as small as $0.52$ mrad ($5.2\times 10^{-4}$ rad). In a more recent study (Li et al., 2025), the BS was also considered to have deviated slightly from the ideal configuration $\pi/4$, with the deviation characterized by $\epsilon\propto 10^{-3}$. Generally, the non-linearity coefficient of the cross-Kerr $g$ is small (Dong et al., 2008). However, recent studies have proposed methods to engineer giant cross-Kerr nonlinearities (Wang et al., 2001; Qian et al., 2016; Kounalakis et al., 2018; van Doai et al., 2019; Doai, 2019; Mu et al., 2021; Huang et al., 2024), with a reported maximum value of $g=0.35$ (Hoi et al., 2013). Reference (Du et al., 2024) explored the implementation of various quantum logic gates based on enhanced cross-Kerr nonlinearity. These findings suggest that the weak value $\langle n_{a}\rangle_{w}$ can be significantly amplified by WVA (Feizpour et al., 2011; Hallaji et al., 2017), making it feasible to achieve the condition $\frac{g}{2}\left(1+\frac{1}{\epsilon}\right)\gg 1$ with appropriate control techniques. This result indicates that the preparation of a new family of quantum states using our proposed scheme is experimentally accessible. In the next section, we present examples that illustrate the preparation of nG states and the enhancement of non-Gaussianity for different input signal states $|\phi\rangle$.
III Non-Gaussian sates generation
III.1 photon-added coherent sate
According to Eq. (8), if the input signal state is taken to be a coherent state, i.e., $|\phi\rangle=|\beta\rangle=D\left(\beta\right)|0\rangle$, where the displacement operator is defined as $D(\beta)=\exp\left(\beta b^{\dagger}-\beta^{\ast}b\right)$, then the normalized final state of the signal beam, after the postselected weak measurement implemented via our protocol, becomes
$|\Phi^{\prime}\rangle$ $=\mathcal{N}\left[|\beta\rangle-ig\langle n_{a}\rangle_{w}b^{\dagger}b|\beta\rangle\right]=\mathcal{N}\left[|\beta\rangle-\kappa|\phi_{1}\rangle\right],$ (11)
where $\kappa=i\beta g\langle n_{a}\rangle_{w}$, $|\phi_{1}\rangle=b^{\dagger}|\beta\rangle$, and the normalization constant $\mathcal{N}$ is given by
$\mathcal{N}$ $=\left[1+|\kappa|^{2}(1+|\beta|^{2})-2Re[\kappa^{\ast}\beta]\right]^{-\frac{1}{2}}.$ (12)
Here, $Re[\mathcal{C}]$ denotes the real part of the complex number $\mathcal{C}$. Furthermore, using the commutation relation $[b^{\dagger},D(\beta)]=\beta^{\ast}D(\beta)$, the output signal state $|\Phi^{\prime}\rangle$ can be rewritten as
$|\Phi^{\prime}\rangle=\mathcal{N}\left[\left(1-\kappa\beta^{\ast}\right)|\beta\rangle-\kappa|\beta,1\rangle\right],$ (13)
where $|\beta,1\rangle=D(\beta)|1\rangle$ denotes the displaced single-photon state. This expression shows that state $|\Phi^{\prime}\rangle$ lies in the subspace spanned by two orthogonal states: the coherent state and the displaced single-photon state. This orthogonality may be useful for encoding CV qubits.
Figure 2: Fidelity functions between the signal output state $|\Phi^{\prime}\rangle$ and the ideal SPAC state (solid curve), as well as the coherent state $|\phi\rangle$ (dashed curve). Here, $\epsilon=0.001$ and $g=0.01$.
We observe that the state $|\Phi^{\prime}\rangle$ is a superposition of a coherent state $|\beta\rangle$ and a SPAC state $|\phi_{1}\rangle=b^{\dagger}|\beta\rangle$. When the weak value $\langle n_{a}\rangle_{w}$ is large, the contribution of the SPAC component $|\phi_{1}\rangle$ dominates over the coherent component. In general, the generated state can be regarded as a typical example of a hybrid coherent state of the form $\sqrt{\varepsilon}|\beta\rangle+\sqrt{1-\varepsilon}|\phi_{1}\rangle$, with $0\leq\varepsilon\leq 1$ (Turek et al., 2023). Such states exhibit enhanced nonclassical features compared to the SPAC state alone because of their coherent-state admixture in specific parameter regimes.
To quantify how closely the generated state $|\Phi^{\prime}\rangle$ resembles either the coherent state $|\beta\rangle$ or the ideal SPAC state $|\phi_{1}\rangle$, we computed the fidelity:
$F_{0}=|\langle\beta|\Phi^{\prime}\rangle|^{2},\ \ F_{1}=|\langle\phi_{1}|\Phi^{\prime}\rangle|^{2}.$ (14)
Here, $F_{0}$ and $F_{1}$ measure the overlap between the coherent and SPAC states, respectively. A fidelity value that is close to unity indicates that the generated state closely approximates the corresponding target state.
In Fig. 2, we plot $F_{0}$ and $F_{1}$ as functions of the coherent state amplitude $\beta$. Here, we fixed the parameters as $\epsilon=0.001$ and $g=0.01$. As seen, for small values of $\beta$, the state is predominantly coherent, with $F_{0}$ near unity. As $\beta$ increases, $F_{1}$ increases, indicating a closer resemblance to the SPAC state. For a sufficiently large $\beta$, the output state again approaches a coherent state regardless of the other parameter values.
To further characterize the output state, we examine its Wigner function, which provides a complete description of the phase-space distribution of the state. For a quantum state with density operator $\rho$, the Wigner function is obtained via the Fourier transform of the characteristic function $C_{w}(\lambda)=Tr\left[\rho D(\lambda)\right]$. In this study, we used the Python package QuTiP (Johansson et al., 2012, 2013) to numerically compute the Wigner functions of the relevant quantum states.
Figure 3: Wigner function of the output state of the generated signal $|\Phi^{\prime}\rangle$. (a) $\beta=0$; (b) $\beta=0.5$; (c) $\beta=1$. Other parameters are the same as those used in Fig. 2.
To visualize the evolution of the phase-space distribution, Fig.3 shows the Wigner functions of the output state $|\Phi^{\prime}\rangle$ for different values of $\beta$. We fixed the parameters as $\epsilon=0.001$ and $g=0.01$, and considered $\beta=0$, $0.5$, and $1$. Figure3(a) corresponds to the input coherent state, exhibiting the expected Gaussian distribution displaced in the phase space. In Fig.3(b), for $\beta=0.5$, the Wigner function of the output state shows significant negativity near the origin, along with squeezing along the $x$-quadrature. This indicates strong nonclassicality and a marked deviation from the initial Gaussian character. In Fig.3(c), as $\beta$ increases further, the negativity gradually diminishes, and the Wigner function approaches a Gaussian profile again (not shown). This trend is consistent with the behavior of the fidelity functions in the large-$\beta$ regime [see Fig. 2]. These numerical results demonstrate that our protocol effectively transforms the initially Gaussian coherent state into a nG state with tunable nonclassicality, depending on the input amplitude $\beta$.
Furthermore, as shown in Fig. 4, the generated state $|\Phi^{\prime}\rangle$ exhibits a higher single-photon component than a coherent state, although this enhancement is not as strong as that of the ideal SPAC state. Additionally, the state demonstrates a pronounced sub-Poissonian photon-number distribution, as evidenced by its photon-number variance remaining smaller than the square root of the mean photon number, i.e., $\triangle n<\sqrt{\langle n\rangle}$.
Because the SPAC state lacks a vacuum component and exhibits a high probability of single-photon detection, it has found various applications in quantum information processing, including quantum key distribution (QKD) (Wang et al., 2014). In practical QKD implementations, sources with sub-Poissonian photon-number statistics and a dominant single-photon component are highly desirable. Based on the above analysis, our generated state $|\Phi^{\prime}\rangle$ satisfies both of these criteria, making it a promising candidate for QKD applications. Therefore, $|\Phi^{\prime}\rangle$ could offer improved performance compared to other commonly used sources, such as weak coherent states (Zhao et al., 2006; Rosenberg et al., 2007) and heralded single-photon sources (Wang et al., 2008, 2007).
Figure 4: Second-order correlation function $g^{(2)}(0)$. The blue solid curve corresponds to the generated signal output state $|\Phi^{\prime}\rangle$, the red and green dashed curves represent the coherent state and ideal SPAC state cases, respectively. Other parameters are the same as those used in Fig. 2.
III.2 Squeezed photon number state
In the previous subsection, we considered a coherent state as the signal input and demonstrated that the resulting output states exhibited nG characteristics. In this subsection, we consider another typical example, a single-mode SV state. The SV state is defined as $|r\rangle=S(r)|0\rangle$, where the squeezing operator is given by $S(r)=\exp\left[\frac{1}{2}(r^{\ast}b^{2}-rb^{\dagger 2})\right]$, and $r$ is the squeezing parameter. For $r>0$, the SV state is squeezed in the position quadrature $x=\frac{1}{\sqrt{2}}(b+b^{\dagger})$ and anti-squeezed in the momentum quadrature $p=\frac{-i}{\sqrt{2}}(b-b^{\dagger})$; the reverse holds for $r<0$. The squeezed vacuum state is a fundamental resource in quantum optics (Braunstein, 2005), with substantial interest owing to its distinctive properties and wide-ranging applications in quantum sensing (Polzik et al., 1992; et al., 2011), quantum communication (Braunstein and van Loock, 2005), and others (see, e.g., Refs. (Vahlbruch et al., 2016; Andersen et al., 2016; Schnabel, 2017)). In addition, the photon-added and photon-subtracted SV states have been studied extensively (Parigi et al., 2007). In general, generating multiphoton-added or multiphoton-subtracted SV states suffers from very low success probabilities due to measurement imperfections and operational inefficiencies. From the implementation perspective, photon subtraction is typically easier than photon addition. Therefore, the preparation of multiphoton-added SV states requires advanced techniques and higher efficiency. Herein, we propose a simple method to generate two-photon-added squeezed vacuum (TPASV) states without requiring actual photon-addition operations. In the schematic of our setup, if the initial signal (pointer) state is prepared as an SV state, the final signal output state after the measurement process can be expressed as
$|\varPsi\rangle$ $=\chi S(r)\left[|0\rangle+\mu|2\rangle\right]$ (15)
where the normalization coefficient $\chi$ equals to $\chi=(1+\mu^{2})^{-1/2}$ with
$\mu=\frac{i\frac{g}{\sqrt{2}}\left(1+\frac{1}{\epsilon}\right)sinh\left(r\right)cosh\left(r\right)}{1-i\frac{g}{2}\left(1+\frac{1}{\epsilon}\right)sinh^{2}\left(r\right)}.$ (16)
The signal output state is a superposition of a SV state and a squeezed two-photon state. We can express $|\Psi\rangle$ in a more compact form as
$|\Psi\rangle=\eta\left[|r\rangle-ig\langle n_{a}\rangle_{w}|\psi_{b^{\dagger 2}}\rangle\right],$ (17)
where
$|\psi_{b^{\dagger 2}}\rangle=\mathcal{N}_{a^{\dagger 2}}S\left(r\right)\left[|0\rangle-\sqrt{2}\left(\tanh r\right)^{-1}|2\rangle\right]$ (18)
is the TPASV state with normalization constant $\mathcal{N}_{a^{\dagger 2}}=[1+2(\tanh r)^{-2}]^{-1/2}$. The normalization constant $\eta$ for the state $|\Psi\rangle$ is given by
$\eta^{-2}=1+g^{2}\langle n_{a}\rangle_{w}^{2}\ sinh^{4}(r)+\frac{g^{2}}{2}\langle n_{a}\rangle_{w}^{2}\ sinh^{2}(2r).$ (19)
Figure 5: Wigner functions (contour plots) and quadrature distributions for the generated state $|\Psi\rangle$ with different squeezing parameters $r$. (a) and (c) show the Wigner functions of the signal input SV states with squeezing parameters $r=-0.7$ and $r=0.2$, respectively. The corresponding Wigner functions of the output signal states $|\Psi\rangle$ are shown in (b) and (d). Other parameters are the same as those used in Fig. 2.
The TPASV state is generated by successively adding two photons to the SV state. Interestingly, this process is also equivalent to subtracting one photon from the SV state, and subsequently adding one photon. That is,
$b^{\dagger 2}S\left(r\right)|0\rangle=b^{\dagger}bS\left(r\right)|0\rangle.$ (20)
This relation can be verified using the following operator identities:
$b^{\dagger}S\left(r\right)|0\rangle$ $=\cosh rS\left(r\right)|1\rangle,$ (21a)
$bS\left(r\right)|0\rangle$ $=-\sinh rS\left(r\right)|1\rangle,$ (21b)
along with the unitary transformations:
$S^{\dagger}\left(r\right)bS\left(r\right)$ $=b\cosh r-b^{\dagger}\sinh r,$ (22a)
$S^{\dagger}\left(r\right)b^{\dagger}S\left(r\right)$ $=b^{\dagger}\cosh r-b\sinh r.$ (22b)
From Eqs. (21a) and (21b), we see that the squeezed single-photon state $S(r)|1\rangle$ can be obtained by adding or subtracting a single photon to or from an SV state. As confirmed in previous studies (Lund et al., 2004; Ourjoumtsev et al., 2006), the SV state and squeezed single-photon state are good approximations of small-amplitude even and odd SC states, respectively. The SC states are defined as
$|SCS_{\pm}(\alpha)\rangle=\mathcal{N}_{\pm}\left(|\alpha\rangle\pm|-\alpha\rangle\right),$ (23)
where the normalization constants are given $\mathcal{N}_{\pm}=(2\pm 2e^{-2\alpha^{2}})^{-1/2}$. When expressed on the Fock basis, the even SC state contains only even photon-number components, whereas the odd SC state consists only of odd-number components.
In our scheme, the second term in the signal output state is $|\psi_{b^{\dagger 2}}\rangle$ [see Eq.(18)], which is proportional to the weak value $\langle n_{a}\rangle_{w}$. As discussed previously, when $\langle n_{a}\rangle_{w}$ has a large value, the second term becomes dominant in the output signal state [see Eq.(17)], enabling us to effectively obtain the TPASV state without performing actual photon-addition operations.
As noted above, the generated state $|\Psi\rangle$ takes the form $c_{0}|0\rangle+c_{2}|2\rangle$, with $c_{0},c_{2}\in\mathbb{C}$. That is, it is a squeezed superposition of the vacuum and two-photon Fock states. Such a state provides an excellent approximation of a squeezed even SC state with large amplitude $\alpha$ (Marek et al., 2008), which is defined as
$|\Psi_{sscs}\rangle=\mathcal{N}_{sscs}S\left(r^{\prime}\right)\left[|\alpha\rangle+|-\alpha\rangle\right],$ (24)
where the normalization constant is $\mathcal{N}_{sscs}=\left(2+2e^{-2\alpha^{2}}\right)^{-1/2}$. This approximation is valid because the squeezing operator $S(r^{\prime})$, with an appropriately chosen squeezing parameter $r^{\prime}$, effectively pulls apart the two peaks of the cat state. To explain this effect more clearly, in Fig. 5, we plotted the Winger functions of our input SV and the corresponding signal output state $|\Psi\rangle$ for different squeezing parameter $r$. Using our protocol, the input Gaussian-type SV state transformed to the nG state characterized by Winger negativity, and squeezing occurred in both $x$ and $p$ quadratures [ see Figs. 5 (b) and (d) ]. The interference fringes and quadrature distribution features of the state $|\Psi\rangle$ showed in Fig. 5 is very similar to the squeezed even SC state which is the basis of GKP states (Gottesman et al., 2001; Hastrup and Andersen, 2022). It also indicates that the two peaks of the generated state are separated along $x$ or $p$ directions depending on the value of the squeezing parameter $r$. In the following, we further analyze the similarity between our generated state $|\Psi\rangle$ and the ideal large-amplitude squeezed even SC state using fidelity and phase-space distributions. In the following, we assume that both squeezing parameters $r$ and $r^{\prime}$ are real.
The fidelity $F_{2}$ between the generated state $|\Psi\rangle$ and the ideal squeezed even SC state $|\Psi_{sscs}\rangle$ is given by
$F_{2}$ $=|\langle\Psi_{sscs}|\Psi\rangle|^{2}$
$=\frac{4\mu e^{\frac{-2\alpha^{2}}{1+\mu^{2}}}}{1+\mu^{2}}\left|\chi\mathcal{N}_{sscs}\left(\sqrt{2}+\mu^{*}\frac{1-\nu^{4}+4\nu^{2}\alpha^{2}}{1+\nu^{2}}\right)\right|^{2},$ (25)
where $\nu=e^{-(r-r^{\prime})}$.
In Fig. 6, we present the optimal fidelity between $|\Psi\rangle$ and $|\Psi_{sscs}\rangle$ as a function of the coherent amplitude $\alpha$ for different squeezing parameter $r$. In these plots, for each value of $\alpha$, we optimize the squeezing parameter $r^{\prime}$ of the ideal squeezed even SC state to maximize its overlap with the generated state $|\Psi\rangle$. The cases $r=-0.7$ and $r=-2$ are shown in panels (a) and (b), respectively. The corresponding optimized values of $r^{\prime}$ as a function of $\alpha$ are plotted in panels (c) and (d), respectively.
In Figs.6(a) and 6(b), the blue solid curves represent the optimal fidelity between $|\Psi\rangle$ and $|\Psi_{sscs}\rangle$, while the green dashed curves show the fidelity between $|\Psi\rangle$ and the ideal even SC state (without squeezing). As shown in Fig.6(a) for $r=-0.7$, the optimal fidelity reaches as high as $F_{2}=0.97$ at $\alpha=1.91$ and $r^{\prime}=-0.14$, indicating a high degree of similarity between the generated state and the ideal squeezed even SC state. As shown in Fig. 6(b), when $r=-2$, the output signal state $|\Psi\rangle$ provides an excellent approximation to the squeezed even SC state, achieving a very high fidelity of $F_{2}=0.994$ at $\alpha=2.07$ and $r^{\prime}=-0.192$. This result indicates that, in this case, the output signal state $|\Psi\rangle$ is extremely well approximated by the ideal squeezed even SC state.
Figure 6: (a) and (b) show the optimal fidelities between the output signal state $|\Psi\rangle$ and the target squeezed even SC state (solid curves), as well as the ideal even SC state (dashed curves), for different system parameters. (c) and (d) show the corresponding squeezing parameter $r^{\prime}$ of the target squeezed even SC state for which the fidelities in (a) and (b) are optimized. The squeezing parameters of the initial SV states are $r=-0.7$ and $r=-2$ for (a) and (b), respectively. The other parameters are the same as those used in Fig. 2.
While fidelity provides a useful quantitative measure of the similarity between the generated state and the target state, it is also informative to examine the Wigner functions, as each quantum state exhibits unique features in its phase-space distribution. To this end, in Fig. 7, we present the Wigner functions of the output signal state $|\Psi\rangle$ and compare them with those of the ideal squeezed even SC state $|\Psi_{sscs}\rangle$.
Figure 7(a) shows the Wigner function of $|\Psi\rangle$ for $r=-0.7$, while Fig.7(b) displays the Wigner function of the squeezed even SC state with $r^{\prime}=-0.14$ and $\alpha=2$. For these parameters, the fidelity between the two states is approximately $F_{2}\approx 0.97$. It also further confirmed the high similarity of our generated output signal state $|\Psi\rangle$ to the $|\Psi_{sscs}\rangle$ with large amplitude $\alpha$. Fig.7 also illustrates the effects of our protocol on the initially prepared Gaussian SV state, which transforms it into a nG state.
Figure 7: (a) Wigner function of the generated state $|\Psi\rangle$ with $r=-0.7$. (b) Wigner function of the ideal squeezed even SC state with $r^{\prime}=-0.14$ and $\alpha=2$. Other parameters are the same as those used in Fig. 2. For these parameters, the fidelity calculated using the Wigner functions of the generated and ideal squeezed even SC states is $F_{2}\approx 0.97$.
As mentioned in the Introduction, the generation of SC states with high fidelity ($F\geq 0.99$) while maintaining large coherent amplitudes ($\alpha\geq 2$) is a crucial task. Although our scheme does not directly generate large-amplitude SC states, the output signal state achieves exceptionally high fidelity with a squeezed even SC state possessing a large amplitude $\alpha$. Therefore, our approach could be applied to problems where large-amplitude SC states are required, offering potentially better performance.
IV Enlargement of Schrödinger cat states
In the preceding discussion, we demonstrated the usefulness of our scheme in generating nG states from Gaussian input states. We now explore the advantages of our protocol for amplifying SC states. In particular, we investigate the possibility of enhancing the non-Gaussianity of nG states by taking SC states as an example.
If the input signal state is one of the SC states defined in Eq. (23), the output signal state produced by our scheme takes the following form:
$|\Psi_{\pm}\rangle$ $=\mathcal{N_{\pm}^{\prime}}\left[\left(1-i\kappa^{\prime}b^{\dagger}b\right)|\alpha\rangle\pm\left(1+i\kappa^{\prime}b^{\dagger}b\right)|-\alpha\rangle\right]$
$=\mathcal{N_{\pm}^{\prime}}\left[|\alpha\rangle\pm|-\alpha\rangle-i\kappa^{\prime}b^{\dagger}b(|\alpha\rangle\pm|-\alpha\rangle)\right]$ (26)
where $\kappa^{\prime}=g\langle n\rangle_{w}$ and $\mathcal{N}_{\pm}^{\prime}$ is the normalization constant, given by
$\mathcal{N_{\pm}^{\prime}}=\frac{1}{\sqrt{2}}\left[\left(\kappa^{\prime 2}\alpha{}^{4}+1\right)\left(1\pm e^{-2\alpha^{2}}\right)+\kappa^{\prime 2}\alpha{}^{2}\left(1\mp e^{-2\alpha^{2}}\right)\right]^{-\frac{1}{2}}.$ (27)
As seen from Eq. (26), the output state is a superposition of the initial SC state and photon-number-weighted SC state $b^{\dagger}b|SCS_{\pm}\rangle$. Since the weak value $\langle n\rangle_{w}$ can become large in our protocol, the second term can dominate, effectively transforming the initial SC state into $b^{\dagger}b|SCS_{\pm}\rangle$ via the WVA.
Although the output state $|\Psi_{\pm}\rangle$ differs slightly in form from an ideal SC state, the deviation is relatively small. To quantitatively evaluate the similarity, we computed the fidelity between $|\Psi_{\pm}\rangle$ and an ideal SC state with a potentially different coherent amplitude $\alpha^{\prime}$, defined as
$F_{\pm}$ $=|\langle SCS_{\pm}(\alpha^{\prime})|\Psi_{\pm}\rangle|^{2}$
$=4\left|\mathcal{N}_{\pm}\mathcal{N}_{\pm}^{\prime}\left[\left(1-i\kappa^{\prime}\alpha^{\prime}\alpha\right)h_{-}\pm\left(1+i\kappa^{\prime}\alpha^{\prime}\alpha\right)h_{-}\right]\right|^{2},$ (28)
where $h_{\pm}=\exp\left[-\frac{1}{2}|\alpha^{\prime}\pm\alpha|^{2}\right]$. In the above calculations, we assume that both $\alpha$ and $\alpha^{\prime}$ are real.
Figure 8: (a) and (c) show the optimal fidelities between the output signal states $|\Psi_{\pm}\rangle$ and the ideal SC states $|SCS_{\pm}(\alpha^{\prime})\rangle$, plotted as a function of the coherent state amplitude $\alpha$. (b) and (d) show the corresponding enlargement coefficients $\eta_{-}=\alpha^{\prime}/\alpha$ and $\eta_{+}=\alpha^{\prime}/\alpha$, for the odd and even SC states, respectively, for which the fidelities in (a) and (c) are optimized. The other parameters are the same as those used in Fig. 2.
In Fig. 8, we show the optimal fidelity functions $F_{+}$ and $F_{-}$ between our generated output states $|\Psi_{+}\rangle$ and $|\Psi_{-}\rangle$ and the ideal even and odd SC states, respectively, as functions of coherent state amplitude $\alpha$. The corresponding optimized amplitudes $\alpha^{\prime}$ of the ideal SC states are used for the maximal overlap. As shown in Fig. 8(a), the generated state $|\Psi_{-}\rangle$ provides an excellent approximation to the ideal odd SC state, with fidelity $F_{-}\geq 0.965$ across a wide range of amplitudes, $\alpha$. Notably, in the regions $0.80<\alpha<$ 0.8 and $\alpha\geq 2.45$, the fidelity exceeds 0.99. For the even SC state input case [Fig. 8(c)], the generated state $|\Psi_{+}\rangle$ initially shows noticeable deviation from the ideal even SC state in the small-amplitude regime ($0.290<\alpha<$ 0.29). However, as $\alpha$ increases beyond 0.29, the fidelity $F_{+}$ improves significantly, surpassing $0.9$ for all $\alpha>1.24$ regions. Thus, for $\alpha>1.24$, $|\Psi_{+}\rangle$ also can be considered an excellent approximation to the ideal even SC state. Moreover, in high-fidelity regions, we observed that the optimized coherent state amplitudes $\alpha^{\prime}$ of the ideal SC states are consistently larger than the corresponding values of $\alpha$ for our generated states $|\Psi_{\pm}\rangle$ [see Figs. 8(b) and 8(d)]. This indicates that our protocol effectively enlarges the SC states. This enlargement effect can be further confirmed by comparing the Wigner functions of the input ideal SC states and the generated output states $|\Psi_{\pm}\rangle$.
Figure 9: Wigner functions of odd SC states. The first row presents the Wigner functions of the input odd SC states for different coherent amplitudes $\alpha$. The second row shows the corresponding Wigner functions of the generated states $|\Psi_{-}\rangle$ for optimized amplitudes $\alpha^{\prime}$, chosen to maximize the fidelity with the ideal odd SC state $|SCS_{-}(\alpha)\rangle$. Other parameters are the same as those used in Fig. 2.
In the small-amplitude regime ($\alpha\leq 1.2$), the odd SC state is well approximated by the squeezed single-photon state $S(r)|1\rangle$, with a fidelity exceeding $0.99$ (Zavatta et al., 2004; Ourjoumtsev et al., 2006). The corresponding optimized squeezing parameter that maximizes the fidelity for a given $\alpha$ is given by $r_{opt}(\alpha)=\ln\left(\sqrt{\frac{2\alpha^{2}}{3}+\frac{1}{3}\sqrt{9+4\alpha^{2}}}\right)$. However, for larger amplitudes ($\alpha\geq 2$), the fidelity rapidly decreased below $87.8\%$. In contrast, we find that our protocol can generate odd SC states with high fidelity, even for large amplitudes $\alpha\geq 2$. In Fig. 9, we plot the Wigner functions of the odd SC input states for different values of $\alpha$, along with the corresponding output states $|\Psi_{-}\rangle$, using the optimized amplitude $\alpha^{\prime}$ which maximizes the fidelity between ||||$\Psi_{-}\rangle$ and the ideal odd SC target state $|SCS_{-}(\alpha)\rangle$. As shown, for each input odd SC state, our protocol yields an excellent approximation to a larger-amplitude odd SC state encoded in the output state $|\Psi_{-}\rangle$. For example, when the input amplitudes are $\alpha=0.5$, $1.52$, and $2.0$, the corresponding enlarged SC states, characterized by the output amplitudes $\alpha^{\prime}=0.85$, $2.00$, and $2.41$, are obtained with high fidelities: $F_{-}=0.999$, $0.968$, and $0.982$, respectively. The amplitude enlargement ratio $\eta_{-}=\alpha^{\prime}/\alpha$ is plotted in Fig. 8(b). The curve shows that $\eta_{-}>1$ across a wide range of $\alpha$, indicating that our protocol effectively amplifies the amplitude of a given odd SC state while maintaining high fidelity.
Figure 10: Wigner functions of even SC states. The first row presents the Wigner functions of the input even SC states for different coherent amplitudes $\alpha$. The second row shows the corresponding Wigner functions of the generated states $|\Psi_{+}\rangle$ for optimized amplitudes $\alpha^{\prime}$, chosen to maximize the fidelity with the ideal even SC state $|SCS_{+}(\alpha)\rangle$. Other parameters are the same as those used in Fig. 2.
The amplitude-enlarging advantage of our protocol for SC states can also be confirmed by using the even SC state as an example. Similar to the odd SC case, our protocol can generate large-amplitude even SC states with high fidelity when a small-amplitude even SC state is used as input. As shown in Fig. 10, when the input even SC states have amplitudes $\alpha=1.5$, $2.0$, and $3.0$, the corresponding enlarged even SC states with amplitudes $\alpha^{\prime}=2.0$, $2.41$, and $3.3$ are obtained, with fidelities $F_{+}=0.950$, $0.981$, and $0.995$, respectively. These enlarged states are encoded in the output state $|\Psi_{+}\rangle$, generated by our protocol for a given ideal small-amplitude input, even SC state. As shown in Fig. 8(d), the amplitude enlargement ratio $\eta_{+}=\alpha^{\prime}/\alpha$ for even SC states remains greater than one for all tested values of $\alpha$, further confirming that our protocol exhibits a robust amplitude enhancing effect for even SC states. Moreover, the even SC state closely resembles a Gaussian squeezed vacuum state when $\alpha\lessapprox 0.75$, and in that regime, it serves as a high-fidelity ($F=0.99$) approximation to the SV state (Brańczyk and Ralph, 2008). Therefore, for even SC inputs with $\alpha\lessapprox 0.75$, the output behavior mirrors the results discussed in Sec. III.2, using the optimized squeezing parameter $r_{opt}(\alpha)=\ln\left(\sqrt{2\alpha^{2}+\sqrt{1+4\alpha^{2}}}\right)$. Although the even SC state is approximately Gaussian in this small-amplitude regime, the corresponding output state generated by our protocol displays strong nG features, as indicated by the appearance of Wigner negativity [see Figs. 5(c) and 5(d)]. The amplitude enlargement effect of our protocol can also be directly confirmed by comparing the average photon numbers of the input and output states. Large-amplitude SC states naturally contain more photons than small-amplitude states. In the output states $|\Psi_{\pm}\rangle$, the first term corresponds to the input SC state, whereas the second term, $b^{\dagger}b|SCS_{\pm}\rangle$, contributes significantly to the photon number. As a result, the average photon number in the output states $|\Psi_{\pm}\rangle$ is always larger than that in the corresponding ideal SC inputs, while still maintaining high fidelity within the discussed parameter regimes.
It should be emphasized that the amplitude-enlarging feature of our protocol for SC states remains valid across a wide range of coherent amplitudes $\alpha$, with the corresponding fidelity $F_{-}$ ($F_{+}$) exceeding 0.99 for $\alpha\gtrsim 2.41$. In these regimes, the enlargement coefficients $\eta_{-}$ ($\eta_{+}$) are also consistently greater than one. The above results indicate that even though we did not directly prepare the exact large-amplitude SC states, the equivalent SC states $|SC_{\pm}(\alpha^{\prime})\rangle$ were encoded in our signal output states $|\Psi_{\pm}\rangle$. Consequently, the two coherent components |±$|\pm\alpha^{\prime}\rangle$ of the equivalent SC states $|SC_{\pm}(\alpha^{\prime})\rangle$ become approximately orthogonal when $\alpha^{\prime}\geq 2$, making such states valuable resources for exploring macroscopic quantum superpositions and quantum information applications.
As discussed in the Introduction, optical SC states can be generated via cross-Kerr interactions characterized by third-order nonlinear susceptibility $\chi^{(3)}$ (Gerry, 1999; Gerry and Knight, 2004). However, in practice, generating large-amplitude SC states under realistic conditions is extremely challenging due to the typically small and attenuated values of $\chi^{(3)}$. As analyzed above, by using the WVA technique to effectively enhance single-photon level nonlinearities, our protocol enables the generation of large-amplitude SC states ($\alpha\geq 2$) with high fidelity, even in the presence of weak cross-Kerr interactions.
In recent years, several methods have been explored for enlarging SC states (Lund et al., 2004; Sychev et al., 2017; Hacker et al., 2019; Li et al., 2024). Among these, the synthesis-based method for enlarging even SC states, as described in Refs. (Lund et al., 2004; Sychev et al., 2017) requires the input to be in a small odd SC state. However, this approach is constrained by the amplitude of the initial state and is resource-intensive, making the generation of large-amplitude SC states experimentally demanding. In Ref. (Hacker et al., 2019), SC states with amplitudes up to $\alpha=1.4$ were generated through direct interaction between a coherent light pulse and a single-sided cavity containing a three-level atom. Although promising, this method is limited in scalability to larger amplitudes, and is susceptible to decoherence when strong light fields are used.
In contrast, our protocol provides a simple and experimentally feasible alternative for generating large-amplitude SC states with high fidelity across a broad range of amplitudes $\alpha$, without requiring explicit photon addition or subtraction operations. However, we note that the success probability of our scheme is not high and is proportional to the square of the beam splitter deviation from the ideal $50:50$ configuration, i.e., $|\epsilon|^{2}$.
V Conclusions
In this study, we propose a feasible protocol for the preparation of the nG quantum state through weak measurement, implementing a weak cross-Kerr interaction characterized by third-order nonlinear susceptibility $\chi^{(3)}$. A single photon was injected into the interference setup through the idler input port. Upon the detection of a single photon at the dark port of the idler beam, the Gaussian input state at the signal port is transformed into a nG state. This transformation is achieved through the amplification of the single-photon nonlinearity, which is enabled by the WVA technique. As specific examples of the protocol, we demonstrate the generation of various nG states, including SPAC states, displaced (squeezed) single- and two-photon states, and TPASV states, all of which have high fidelity.
Furthermore, this protocol enables the generation of a continuous range of nG states with intermediate properties that depend on the initial state of the signal. We also show that the protocol can be employed to enhance the non-Gaussianity of nG states. We find that, using SC states as a test case, the Wigner function negativity, key indicator of non-Gaussianity, significantly increased after passing through our setup. In particular, the protocol also effectively amplifies small SC states into larger ones with coherent amplitudes $\alpha\geq 2$ while maintaining high fidelity ($F>0.99$), particularly in the case of odd SC states.
Throughout the process, the transformation of the Gaussian input into a nG state is primarily owing to an effective sequence of single-photon subtraction followed by addition in the signal mode, facilitated by the weak cross-Kerr interaction and WVA. Remarkably, photon addition and subtraction occur without the need for explicit nG operations. Our results further confirm the utility of postselected weak measurements in quantum state engineering and suggest that the ability to generate a continuous family of nG states may find practical applications in a variety of quantum information processing tasks.
Acknowledgements.
This study was supported by the National Natural Science Foundation of China (No. 12365005).